# Large Language Model Performance Benchmarking on Mobile Platforms: A Thorough Evaluation

### v1

@@ 5.2 Impact of DVFS
In Section 4.1.1, we observed performance degradation as the length of the prompt and generation increased. This issue relates to the system’s ability to handle long contexts and maintain consistent inference performance across multiple rounds of dialogue. Although previous studies, such as [25] and [27], have discussed potential causes like Dynamic Voltage and Frequency Scaling (DVFS) and thermal throttling, they lack detailed analysis on how DVFS specifically impacts inference performance and whether its effects differ across hardware and operating systems. To address this, we analyze real-time variations in CPU frequency and utilization during 20 continuous inference tests on four devices with warm up, covering different operating systems (Android, HyperOS, HarmonyOS) and vendors (MediaTek, Qualcomm, Hisilicon).
@@
Figure 9 illustrates a general decline in performance as the rounds of inference increase. Specifically, on the Snapdragon Gen (on Xiaomi14 Pro), latency increases by up to , dropping from tokens/s to tokens/s. In contrast, Dimensity and Kirin E exhibit only minor latency fluctuations, with a maximum decrease of approximately . This trend aligns with the data presented in Figure 3, which shows that the Xiaomi Pro experiences a more significant performance drop as prompt length increases.
@@
Figure 9 shows the frequency variation on the prime CPU core at different rounds of inference. It reveals that Snapdragon SoCs (e.g.Huawei Matepad11 Pro, Xiaomi14 Pro) exhibit a more aggressive DVFS policy. During subsequent stages of the inference, both prime and performance cores on Snapdragon 8 Gen 3 experience a rapid reduction in frequency. By the th inference round, the frequency is nearly halved compared to its initial value. To investigate the factors influencing DVFS policy, we further compare two Huawei tablets equipped with identical CPU cores but different SoCs—one powered by Snapdragon (Snapdragon 870 on Huawei Matepad11 Pro) and the other by Kirin (Kirin 9000E on Huawei Matepad12.6 Pro). The results indicate that the vendor plays a significant role in shaping the DVFS policy. For instance, although the maximum frequencies of both devices are GHz, the Snapdragon 870 typically keeps the prime core frequency below GHz, while the Kirin 9000E sustains a higher frequency of around GHz during inference.
5.3 Model Preparation Latency
@@
In Section 5.2, we have studied the inference performance with warm-up. However, in real-world applications, users may trigger the model sporadically, leading to scenarios where the model must perform a cold start without a warm-up, known as the cold start. Thus, the time required to generate the first token becomes critical to user experience. The process of generating the first token involves several stages: loading model weights from external storage into memory, preparing memory for variables, and executing prefill and decoding. In this section, we focus on the preparation procedure before prefill and evaluate the latency across various devices. We perform tests on six devices and measure the time spent on preparation with and without warm-up. If the device is tested without warm-up, we reboot it to refresh RAM and cache. Results are present in Figure 10.
@@
For the llama.cpp running on CPUs, we compare the time before prefill of the first token in both cold-start and warm-up scenarios. The Dimensity 9300 (on Vivo Pad3 Pro) delivers the best overall performance, with an average model loading time of approximately ms. The Snapdragon 8 Gen 3 (on Xiaomi14 Pro) ranks second, with its worst-case performance comparable to that of the Dimensity 9300. We recognize that the warm-up process significantly reduces preparation time for llama.cpp, primarily due to the use of mmap, which allows the inference process to directly access model weight files from disk. The mmap function accelerates loading by eliminating the need to copy data between user and kernel space. Additionally, the mapping remains valid for other processes accessing the same file, meaning that after a warm-up, the model weights are already loaded into memory. This allows subsequent inference processes to use the preloaded weights, thereby reducing latency. However, this benefit depends on the memory retaining the loaded weights—if they are overwritten by other files, the warm-up advantage is lost. From Figure 10, we observe that all Huawei devices exhibit unusually long preparation times during cold starts. On Huawei devices with Kirin SoCs, the delays are even more pronounced, with the worst loading times exceeding seconds. Huawei Matepad 11 Pro, despite using a Snapdragon SoC, also experiences extended delays. We identify the underlying causes for these exceptions are related to the mmap function in Harmony OS and we will discuss it in Section 5.4.1.
In contrast, for MLC LLM running on GPUs, no significant improvement is observed after warm-up, with loading times remaining approximately five times longer on Adreno GPUs and seven times longer on Mali GPUs compared to llama.cpp (with warm-up). This is due to mobile system architecture. Although the CPU and GPU on mobile devices share the same physical DRAM, they operate in separate memory spaces and cannot directly access each other’s data[47]. As a result, model weights must be copied from the CPU memory space to the GPU memory space, introducing additional latency. Specifically, MLC LLM uses OpenCL kernels to transfer data between the CPU and GPU via buffers, which contributes to the delay.
Type | Memory(RAM) | Storage(ROM) | L3 Cache |
XiaoMi14 Pro | 4800 MHz(LPDDR5X) | UFS 4.0 | 12MB |
Vivo Pad3 Pro | 4800 MHz(LPDDR5X) | UFS 4.0 | 18MB |
The variations in preparation times across devices are primarily linked to memory bandwidth and UFS speed. For example, the preparation times for Xiaomi14 Pro (Snapdragon Gen ) and Vivo Pad3 Pro (Dimensity ) are comparable on both CPU and GPU. However, the subtle hardware differences still influence performance. Snapdragon 8 Gen 3 shows a slight advantage on the GPU, while the Dimensity performs better on the CPU. As shown in Table 4 and Table 5, the Snapdragon Gen (Adreno 750) benefits from higher real-time GPU memory bandwidth, whereas the Dimensity 9300’s larger L3 cache allows for more efficient reuse of loaded weights after the warm-up. verall, preparation times on both CPUs and GPUs across different vendors are short enough to meet the demands of most real-time applications.
5.4 Inference Engines
In the previous sections, we studies the impact of hardware on LLM inference. However, some performance discrepancies stem from the design and implementation of inference engines rather than inherent hardware capability limitations. In this part, we will focus on the role of inference engines and provide recommendations for framework developers to improve performance.
5.4.1 llama.cpp
@@ 5.4 Inference Engines
The first issue concerns the exceptionally long model preparation time without warm-up on Huawei devices. Through timestamp analysis, we identify that the primary cause is the execution of the mmap function. On the Huawei Matepad 12.6 Pro, mmap execution takes up to seconds, whereas after warm-up, this time is drastically reduced to seconds. This suggests that the delay arises from the memory mapping process rather than loading from UFS. We further verify this by testing without file mapping, which reduces the average preparation time to seconds. Additionally, on the Huawei Matepad 11 Pro with a Snapdragon SoC, the average preparation time is seconds—showing noticeable improvement over other Huawei devices but still significantly slower than the non-Huawei devices. These findings suggest that the issue may be related to mmap behavior on HarmonyOS, possibly exacerbated by Kirin SoCs.
Secondly, we compare different versions of llama.cpp on the Snapdragon 8 Gen 3, Dimensity 9300, and Kirin 9000E in Figure 11 to evaluate how code optimizations affect performance across various SoCs. The results show that new-generation SoCs such as Snapdragon 8 Gen 3 and Dimensity 9300 experience significantly greater performance gains compared to others. Additionally, the latest version of llama.cpp optimizes model weight arrangement to align with INT8 matrix multiplication instructions, resulting in a peak prefill speed of up to 56 tokens/s. This improvement highlights that to better harness the full capabilities of processors, developing support for these advanced instructions is necessary.
5.4.2 MLC LLM
In Section 5.1.3, we explored the factors contributing to the suboptimal performance of MLC LLM on non-Snapdragon SoCs. While we attribute this issue to insufficient hardware unit utilization, it is important for developers to understand what implementation optimizations can address this inefficiency. We observe that MLC LLM supports local execution on devices like the Orange Pi [39], which features a Mali-G610 GPU. Despite the fact that Mali-G’s theoretical computation capability being significantly lower than the Mali-G on the Dimensity 9300, the prefill speed on the Orange Pi is tokens per second—more than twice that of the Mali-G. This discrepancy suggests that theoretical computation capability alone is not sufficient to explain performance differences and points to the importance of optimization. Furthermore, MLC LLM configurations vary widely across target devices. For instance, the loop unroll count is set to on Adreno but for Mali, underscoring the necessity of optimizing operators for specific hardware accelerators. Given the greater heterogeneity of mobile GPUs compared to CPUs, achieving optimal performance requires tailored optimizations for each platform.
6 Implications
In this section, we investigate several aspects that can improve mobile LLM performance.
Memory Bottleneck According to Figure 6, the prefill speeding is times faster than decode. Decoding is primarily constrained by memory bandwidth. Current optimizations, such as adding more cores and using specialized instruction sets, have proven effective in accelerating prefill. However, these approaches offer limited improvements for decoding speed. To address this challenge, more effective solutions are needed. Developers should consider designing algorithms that optimize memory utilization during decoding.
@@ 6 Implications
Hardware Instructions A key aspect of SoC evolution is the continuous enhancement of instruction sets. New machine instructions often increase operand width, resulting in higher throughput. To fully harness the potential of hardware, developers must consider the characteristics of model operators and take advantage of hardware-accelerated instructions. This may also involve adjusting computational precision, such as quantizing activations to -bit to utilize instructions like smmla. In Section 5.1.2, we discuss the improvement brought by Armv8 instructions. For next-generation CPUs based on the Armv9 architecture, advanced vector instructions like SVE2[10]- which enables longer computation vectors-are supported. Furthermore, Arm’s SME instructions[8], specifically designed for matrix multiplication, present additional optimization opportunities. Staying current with these advancements is crucial for developers seeking to maximize inference speed on mobile platforms.
@@
7 Discussion
In this section, we explain the rationale behind our choices in this measurement study.
@@ 7 Discussion
Multimodality Models. Driven by the growing demand for advanced AI assistants, the deployment of large multimodal models (LMMs) on mobile devices is an inevitable trend. However, we excluded LMMs from our study for two main reasons. First, current LMMs may not yet be ready for mobile deployment. For example, MobileAIBench reports that testing Llava-Phi-2[48] on the iPhone 14 resulted in a first token generation of seconds. It remains unclear whether this poor performance is due to hardware constraints or inefficient operator implementation. Second, since LLMs and LMMs share significant similarities in both model architecture and operator implementations, the insights gained from on-device LLM tests can likely be extrapolated to large multimodal models.
8 Concluding Remarks
In this paper, we present a comprehensive measurement study of LLM inference on mobile platforms, offering insights into both user-experience metrics (such as token throughput and energy consumption) and critical hardware and system characteristics (including CPU/GPU utilization, DVFS, and file mapping latency). Our analysis reveals how hardware capabilities and system dynamics impact on-device LLM performance and highlights potential bottlenecks affecting mobile LLM applications. Additionally, we propose potential directions for enhancing on-device LLM inference performance. We hope that this study can provide insights for both the development of on-device LLMs and the design for future mobile system architecture.

### v3

@@ V-C1 Impact of DVFS on CPU
In Section IV-A1, we observed performance degradation as the length of the prompt and generation increased. This issue relates to the system’s ability to handle long contexts and maintain consistent inference performance across multiple rounds of dialogue. We analyze real-time variations in CPU frequency and utilization during 20 continuous inference tests on four devices with warm up, covering different operating systems (Android, HyperOS, HarmonyOS, iPadOS) and vendors (MediaTek, Qualcomm, Hisilicon, Apple). Given the significant performance advantage of the Apple device, we performed additional testing using longer prompts and outputs.
@@
Figure 11 illustrates a general decline in performance as the rounds of inference increase. Although the inference throughput for both prompt lengths is similar, the iPad exhibits a more pronounced performance drop in long-prompt tests compared to short-prompt tests. Moreover, in short-prompt tests, unlike the stable prefill speed, the decode speed on the iPad decreases during the early stages and eventually falls from the best to a level comparable with the second-highest among the other devices. Additionally, on Xiaomi 14 Pro, throughput for both prefill and decode initially rises, then decreases, and finally stabilizes at a relatively steady level, accompanied by minor oscillations around the equilibrium. In contrast, Dimensity (Vivo Pad3 Pro) and Kirin E (Huawei Matepad12.6 Pro) exhibit only minor latency fluctuations, with a maximum decrease of approximately . This trend aligns with the data presented in Figure 3, which shows that Xiaomi Pro experiences a more significant performance drop as prompt length increases.
@@
Figure 11 shows the frequency variation on the prime CPU core at different rounds of inference. CPU frequency is inaccessible on the iPad, resulting in a lack of corresponding measurement data. The results reveal that Snapdragon SoCs (e.g., Huawei Matepad11 Pro, Xiaomi14 Pro) exhibit a more aggressive DVFS policy. During subsequent stages of the inference, both prime and performance cores on Snapdragon 8 Gen 3 experience a rapid reduction in frequency. By the th inference round, the frequency is nearly halved compared to its initial value. To investigate the factors influencing DVFS policy, we further compare two Huawei tablets equipped with identical CPU cores but different SoCs—one powered by Snapdragon (Snapdragon 870 on Huawei Matepad11 Pro) and the other by Kirin (Kirin 9000E on Huawei Matepad12.6 Pro). The results indicate that the vendor plays a significant role in shaping the DVFS policy. For instance, the Snapdragon 870 typically keeps the prime core frequency lower than the Kirin 9000, although they have the same maximum frequency.
Implication 11
CPU frequency critically affects inference performance. The default DVFS strategy tends to be conservative, taking factors such as thermal throttling and battery life into consideration. While this strategy is suited for successive multi-round inferences, the more common case on mobile devices involves intermittent single-round inference (with larger than s interval). Therefore, developers may consider increasing the CPU frequency—or even overclocking—for initial inference to reduce response latency.
V-C2 Impact of CPU Thermal Throttling
As discussed in Section V-C1, only a few consecutive inference runs are sufficient to trigger CPU frequency reduction due to thermal throttling. To further examine the impact of heat accumulation over time, we extend the experiments on Xiaomi and analyze () temperature and performance variation during sustained Inference and () the impact of initial device temperature on performance.
Impact of Heat Accumulation on Sustained Inference: Figure 13 illustrates performance variations of different workloads over consecutive inference runs and the corresponding thermal buildup. Both curves exhibit an initial rapid temperature rise followed by stabilization. During the first seconds, the trends are nearly identical, starting from about and climbing above . In the stable stage, the rate of temperature rise slows, whereas the long-prompt tests are still faster than the short-prompt tests, leading to a higher equilibrium near . After about seconds, the temperature of long-prompt tests fully stabilizes, with minor oscillations around equilibrium likely reflecting stricter thermal regulation.
The performance variation closely follows the temperature trend. During the initial heating stage, throughput fluctuates noticeably. For long-prompt tests, since each inference run takes longer, performance becomes consistent after only – runs. In contrast, short-prompt tests stabilize much later, requiring nearly runs. Across both prefill and decode stages, the performance gap between the two tests is initially narrow, then widens as the system heats up, and finally narrows again and stabilizes at a steady difference of about token/s.
Performance Sensitivity to Starting Temperature: To evaluate performance sensitivity to starting temperature, experiments are conducted under two conditions: a cool-down state (below C) and a warm-up state (above C). These conditions are achieved by first running additional tasks to heat the device and then cooling it for varying durations until the target temperature is reached. All other experimental configurations are largely consistent with Section 1.
Figure 13 shows that overall performance decreases at higher temperatures, with smaller workloads being more sensitive. Although temperature initially rises at a similar rate across different prompt lengths, long-prompt tests require more time per run, causing the CPU frequency to drop rapidly within the first one or two rounds of interference. Consequently, the effect of the initial temperature for long-prompt tests is largely limited to these early rounds, whereas for short-prompt tests, all five consecutive runs occur during the temperature rise stage, allowing lower starting temperatures to delay CPU throttling and sustain higher performance throughout.
V-C3 Impact of DVFS on GPU
We further evaluate the impact of DVFS on GPUs, using llama.cpp since continuous inference with MLC LLM often crashes. Due to framework limitation, only Xiaomi and Apple devices are tested and the results are illustrated in Figure 14. Leveraging faster GPU inference with llama.cpp, experiments are extended to runs with larger workloads (-token prompts, -token outputs). GPU frequency data are available only on the rooted Xiaomi Pad Pro, while iPad reports thermal states via Xcode.
Adreno GPU under Android: The results show that inference speed on both the Xiaomi Pro and Xiaomi Pad Pro remains relatively stable throughout the tests. GPU frequency on Xiaomi Pad Pro is generally fixed at MHz during inference. However, in the first run of some tests, it briefly spikes to MHz before quickly returning to MHz, where it remains stable for subsequent runs. This initial fluctuation is likely caused by the DVFS policy.
Apple GPU under iOS: Unlike the stable performance observed on Adreno GPUs under Android, the iPad exhibits noticeable fluctuations throughout the inference process. For short-prompt tests, the performance degradation is minor and appears near the end of the run. However, for long-prompt tests, the degradation occurs earlier and becomes more pronounced. These results are consistent with the thermal state reported: during most of the 30 inference runs, the device remains in the “nominal” state (within normal thermal limits), but towards the end of long-prompt inferences, the state may shift to “serious”, indicating elevated heat conditions.
Implication 12
Compared with the CPU, whose frequency fluctuates frequently, the GPU maintains a relatively stable frequency during inference, yielding more consistent performance. This stability likely benefits from the GPU’s higher computational efficiency and lower power consumption. These observations suggest that offloading LLM inference to GPUs is a practical and effective strategy for preserving stable real-time responsiveness on mobile devices.
Implication 13
Mobile GPUs may not fully utilize their theoretical computational capacity. For example, the Xiaomi Pad Pro GPU supports frequency levels ranging from MHz to MHz, yet its default operating frequency is MHz, the third-highest level.
Implication 14
The end-to-end inference speed suggests that Apple’s DVFS strategy for the GPU is more conservative than that of Adreno GPUs on Android. Although the Apple GPU offers higher computational capability, it is more prone to performance degradation under sustained workloads due to stricter frequency-scaling policies.
V-D Model Preparation Latency
@@
In Section V-C1, we have studied the inference performance with warm-up. However, in real-world applications, users may trigger the model sporadically, leading to scenarios where the model must perform a cold start without a warm-up, known as the cold start. Thus, the time required to generate the first token becomes critical to user experience. The process of generating the first token involves several stages: loading model weights from external storage into memory, preparing memory for variables, and executing prefill and decoding. In this section, we focus on the preparation procedure before prefill and evaluate the latency across various devices. We perform tests on seven devices and measure the time spent on preparation with and without warm-up. If the device is tested without warm-up, we reboot it to refresh RAM and cache. Results are present in Figure 15.
For the llama.cpp running on CPUs, we compare the time before prefill of the first token in both cold-start and warm-up scenarios. The Xiaomi Pro performs the fastest after warm-up, while its preparation time for cold start is nearly ten times slower than for warm-up. The Vivo Pad Pro delivers the best overall performance for both warm-up and cold start. Additionally, the Apple device exhibits significant fluctuations during cold start, and its warm-up latency is noticeably slower than that of other top-tier devices.
@@
We recognize that the warm-up process significantly reduces preparation time for llama.cpp, primarily due to the use of mmap, which allows the inference process to directly access model weight files from external memory. The mmap function accelerates loading by eliminating the need to copy data between user and kernel space. Additionally, the mapping remains valid for other processes accessing the same file, meaning that after a warm-up, the model weights are already loaded into memory. This allows subsequent inference processes to use the preloaded weights, thereby reducing latency.
@@
From Figure 15, we observe that all Huawei devices exhibit unusually long preparation times during cold starts. On Huawei devices with Kirin SoCs, the delays are even more pronounced, with the worst loading times exceeding seconds. Huawei Matepad 11 Pro, despite using a Snapdragon SoC, also experiences extended delays. We identify the underlying causes for these exceptions are related to the mmap function in Harmony OS. Through timestamp analysis, we identify that the primary cause is the execution of the mmap function. On the Huawei Matepad 12.6 Pro, mmap execution takes up to seconds, whereas after warm-up, this time is drastically reduced to seconds. This suggests that the delay arises from the memory mapping process rather than loading from UFS.
In contrast, for MLC LLM running on GPUs, no significant improvement is observed with warm-up, with loading times remaining approximately nine times longer on Adreno GPUs and five times longer on Mali GPUs compared to llama.cpp (with warm-up). An exception is that the Apple GPU performs slightly better than the CPU when using llama.cpp (with warm-up). This is probably due to the difference between the system memory architectures. For non-Apple devices, although the CPU and GPU on mobile devices share the same physical DRAM, they operate in separate memory spaces and cannot directly access each other’s data[51]. As a result, model weights must be copied from the CPU memory space to the GPU memory space , introducing additional latency. Specifically, MLC LLM uses OpenCL kernels to transfer data between the CPU and GPU via buffers, which results in the delay. In contrast, a major advantage of Apple devices lies in their unified memory architecture [52], which enables shared access between CPU and GPU and thereby achieves much lower transfer latency.
VI Opportunities for Improving Mobile LLM Performance
Building on the results present in the previous sections, this section highlights several key aspects that developers can leverage to significantly enhance mobile LLM performance.
@@ VI Opportunities for Improving Mobile LLM Performance
Hardware Instructions A key aspect of SoC evolution is the continuous enhancement of instruction sets. New machine instructions often increase operand width, resulting in higher throughput. To fully harness the potential of hardware, developers must consider the characteristics of model operators and take advantage of hardware-accelerated instructions. This may also involve adjusting computational precision, such as quantizing activations to -bit to utilize instructions like smmla. In Section V-A2, we have discussed the improvement brought by SVE instructions. For next-generation CPUs based on the Armv9 architecture, advanced vector instructions like SVE2[53], which enables longer computation vectors, are supported. Furthermore, Arm’s SME instructions[54], specifically designed for matrix multiplication, present additional optimization opportunities. Staying current with these advancements is crucial for developers seeking to maximize inference speed on mobile platforms.
Dynamic Resource Allocation Parallel execution of large language models (LLMs) alongside other tasks remains a significant challenge. Optimal resource allocation is inherently dynamic, influenced by real-time workloads and system thermal conditions. A practical approach is to pre-define multiple operating modes that can adapt to varying workloads, such as adjusting the number of threads or offloading computations to idle processors.
@@
Enhancing Memory Access Efficiency To address GPU underutilization, optimization efforts should primarily target memory access efficiency. Potential strategies include restructuring data layouts to improve locality, applying memory coalescing to reduce cache misses, and minimizing reliance on external memory. Furthermore, for long-context inference, partitioning large matrices into smaller chunks can effectively alleviate memory pressure and enhance execution efficiency.
Customized Operators and Kernel Parameter Tuning Customized operator implementations are essential for fully exploiting GPU capabilities and achieving performance breakthroughs. Owing to the significant architectural differences across GPUs, generic operator implementations often fall short in delivering optimal acceleration. Fine-grained parameter tuning of kernel functions, such as adjusting the work size, offers a promising optimization path. For example, an excessively large global work size may introduce severe memory contention, while an overly small size can result in underutilization of compute units and additional overhead.
Dynamic GPU Frequency The computing capability of the GPU is not fully utilized, as its operating frequency is generally fixed at a value lower than the maximum during inference. While this strategy helps maintain stable performance with moderate energy consumption, incorporating dynamic frequency adjustment could achieve a more adaptive balance between performance and energy efficiency.
Hybrid Inference to Enhance User Experience Although CPU performance appears sufficient, as shown in Section V-C2, running inference solely on CPUs leads to rapid temperature increases. Given the strong performance of the NPU and GPU during the prefill stage, a hybrid inference approach—utilizing the NPU or GPU for prefill and the CPU for decoding—is expected to enhance efficiency while delaying temperature rise and improving the overall user experience.
VII Discussion
@@
Multimodality Models. Driven by the growing demand for advanced AI assistants, the deployment of large multimodal models (LMMs) on mobile devices is an inevitable trend. However, we excluded LMMs from our study for two main reasons. First, current LMMs may not yet be ready for mobile deployment. For example, MobileAIBench reports that testing Llava-Phi-2[55] on the iPhone 14 resulted in a first token generation of seconds. It remains unclear whether this poor performance is due to hardware constraints or inefficient operator implementation. Second, since LLMs and LMMs share significant similarities in both model architecture and operator implementations, the insights gained from on-device LLM tests can likely be generalized to large multimodal models.
VIII Concluding Remarks
In this paper, we present a comprehensive measurement study of LLM inference on mobile platforms, offering insights into both user-experience metrics (such as token throughput and energy consumption) and critical hardware and system characteristics (including CPU/GPU utilization, DVFS, and file mapping latency). Our analysis reveals how hardware capabilities and system dynamics impact on-device LLM performance and highlights potential bottlenecks affecting mobile LLM applications. Additionally, we propose potential directions for enhancing on-device LLM inference performance, including customizing operators by tuning kernel parameters to match various GPUs, optimizing data layout to improve memory efficiency, and searching for more adaptive DVFS policies to balance performance and energy. We hope that this study can provide insights for both the development of on-device LLMs and the design for future mobile system architecture.